# Can Few-shot Work in Long-Context? Recycling the Context to Generate Demonstrations

### v1

@@ Abstract
Despite recent advancements in Large Language Models (LLMs), their performance on tasks involving long contexts remains sub-optimal. In-Context Learning (ICL) with few-shot examples may be an appealing solution to enhance LLM performance in this scenario; However, naïvely adding ICL examples with long context introduces challenges, including substantial token overhead added for each few-shot example and context mismatch between the demonstrations and the target query. In this work, we propose to automatically generate few-shot examples for long context QA tasks by recycling contexts. Specifically, given a long input context (1-3k tokens) and a query, we generate additional query-output pairs from the given context as few-shot examples, while introducing the context only once. This ensures that the demonstrations are leveraging the same context as the target query while only adding a small number of tokens to the prompt. We further enhance each demonstration by instructing the model to explicitly identify the relevant paragraphs before the answer, which improves performance while providing fine-grained attribution to the answer source. We apply our method on multiple LLMs and obtain substantial improvements on various QA datasets with long context, especially when the answer lies within the middle of the context. Surprisingly, despite introducing only single-hop ICL examples, LLMs also successfully generalize to multi-hop long-context QA using our approach.
@@ 1 Introduction
We applied DoubleDipper to a variety of LLMs, both commercial (Gemini Pro and Gemini Ultra; Team et al., 2023) and open-source (Llama (Touvron et al., 2023), Mistral (Jiang et al., 2023) and Gemma (Team et al., 2024)), and evaluate it on various QA datasets with long inputs, including common multi-hop QA datasets. Our experiments demonstrate that with only 3 self-generated few-shot examples, DoubleDipper consistently outperforms the baseline on our evaluation set. In addition, for some models, DoubleDipper enhances the robustness to the position of the relevant information within the text. Interestingly, while our few-shot examples focus on single-paragraph answers, DoubleDipper generalizes well to multi-hop QAs where the answer requires information from multiple passages.
@@ 5 Results
Table 3 presents the QA performance of our two variants of DoubleDipper, namely DoubleDipper (Self) and DoubleDipper (PaLM 2) on our evaluation set. We report the performance of the supporting paragraphs prediction in Appendix B. On average, our straightforward method, which automatically generates few-shot examples from the provided input context, outperforms the baseline, often by a significant margin (e.g., +15.2 points for Gemini Pro, +7 points for Mistral Instruct, +6.3 points for Llama2 70B). Comparing the two variants, DoubleDipper (PaLM 2) generally yields better results than DoubleDipper (Self), indicating that higher-quality question-answer (QA) pairs in the few-shot examples enhance performance. Nonetheless, DoubleDipper (Self) already surpasses the baseline by a considerable margin on its own in most benchmarks. Notably, while DoubleDipper produces simple QAs answerable from a single paragraph, it demonstrates strong generalization across diverse QA formats (except SRT), including PIR and MonoRel, which involves True/False questions, and multi-hop QA datasets (HotPotQA, 2Wiki, and MuSique), where reasoning over multiple paragraphs is required.

### v2

@@ Abstract
Despite recent advancements in Large Language Models (LLMs), their performance on tasks involving long contexts remains sub-optimal. In-Context Learning (ICL) with few-shot examples may be an appealing solution to enhance LLM performance in this scenario; However, naïvely adding ICL examples with long context introduces challenges, including substantial token overhead added for each few-shot example and context mismatch between the demonstrations and the target query. In this work, we propose to automatically generate few-shot examples for long context QA tasks by recycling contexts. Specifically, given a long input context (1-3k tokens) and a query, we generate additional query-output pairs from the given context as few-shot examples, while introducing the context only once. This ensures that the demonstrations are leveraging the same context as the target query while only adding a small number of tokens to the prompt. We further enhance each demonstration by instructing the model to explicitly identify the relevant paragraphs before the answer, which improves performance while providing fine-grained attribution to the answer source. We apply our method on multiple LLMs and obtain substantial improvements (+23% on average across models) on various QA datasets with long context, especially when the answer lies within the middle of the context. Surprisingly, despite introducing only single-hop ICL examples, LLMs also successfully generalize to multi-hop long-context QA using our approach.
@@ 1 Introduction
We applied DoubleDipper to a variety of LLMs, both commercial (Gemini Pro and Gemini Ultra; Team et al., 2023) and open-source (Llama (Touvron et al., 2023), Mistral (Jiang et al., 2023) and Gemma (Team et al., 2024)), and evaluate it on various QA datasets with long inputs, including common multi-hop QA datasets. Our experiments demonstrate that with only 3 self-generated few-shot examples, DoubleDipper consistently outperforms the baseline on our evaluation set by 23% on average across models. In addition, for some models, DoubleDipper enhances the robustness to the position of the relevant information within the text. Interestingly, while our few-shot examples focus on single-paragraph answers, DoubleDipper generalizes well to multi-hop QAs and where the answer requires information from multiple passages.
@@ 5 Results
Table 3 presents the QA performance of our two variants of DoubleDipper, namely DoubleDipper (Self) and DoubleDipper (PaLM 2) on our evaluation set. We report the performance of the supporting paragraphs prediction in Appendix B. On average across models and datasets, our straightforward method, which automatically generates few-shot examples from the provided input context, yields an improvement of 12% for DoubleDipper (Self) and 23% for DoubleDipper (PaLM 2), often by a significant margin (e.g., +15.2 points for Gemini Pro, +7 points for Mistral Instruct, +6.3 points for Llama2 70B). Notably, while DoubleDipper produces simple QAs answerable from a single paragraph, it always surpasses the baseline in multi-hop QA datasets (HotPotQA, 2Wiki and MuSique).